OpenStudy (anonymous):

The volume of a cylinder is V = pi r2h. Solve for h (height) in this formula.

jigglypuff314 (jigglypuff314):

V = (pi)(r^2)(h) divide both sides by (pi)(r^2) so h = (V)/((pi)(r^2))